# Privacy-preserving fall detection at the edge using Sony IMX636 event-based vision sensor and Intel Loihi 2 neuromorphic processor

### v1

@@ I Introduction
The world population is rapidly growing with an increase in life expectancy, particularly in developed countries [Bloom_Luca16]. Consequently, elderly care is a global concern which is gaining attention. In particular, falls are becoming a major public health problem for the elderly [Tasnim_Eisuke23]. According to the World Health Organization, falls are the second most common cause of unintentional injury-related deaths worldwide, after road traffic accidents [WHO_21]. Therefore, there is a growing need for fall detection systems in the healthcare industry. A fall detection system aims to trigger an alert after a fall has occurred to emergency caregivers [Wang_etal20]. On the one hand, quick response is critical after a fall and medical outcomes of falls as well as the associated healthcare expenses largely depend on rescue time [Mubashir_etal13, Gharghan_Hashim24]. On the other hand, it has been shown that fall detection systems decrease the fear of falling, which in turn can decrease the likelihood of a fall [Friedman_etal02, Mubashir_etal13].
@@
Fall detection systems use two main types of sensors [Chen_etal17]: body-worn inertial sensors and vision sensors. Accelerometers and gyroscopes embedded in smartphones and smartwatches measure body motion independently of environmental factors (e.g., objects, lighting). However, continuous wear and regular recharging limit their practicality, particularly for elderly people [Alam_etal22]. In contrast, vision sensors like RGB cameras, infrared cameras, depth cameras, and 3D-based methods using camera arrays are non-intrusive for the user and provide continuous passive monitoring without depending on user compliance. They are therefore useful in settings such as nursing homes, hospitals, or assisted living facilities. Combined with deep learning models, vision-based solutions can be highly accurate and adaptable to different environments [Shojaei-Hashemi_etal18]. Yet, privacy concerns remain substantial [Serpanos_etal08, Igual_etal13] as residents often oppose in-room cameras [Xu_etal18]. Studies on fall detection usually lack strategies to effectively ensure data privacy, which prevents their real-life deployment [Igual_etal13].
@@
In this work, we propose a privacy-preserving vision-based solution that consists of the Sony®111Sony is a registered trademark of Sony Group Corporation or its affiliates. Event-based Vision Sensor (EVS) [Finateu_etal20] which only captures per-pixel changes of light intensity in the form of events, combined with the Intel®222Intel is a registered trademark of Intel Corporation or its affiliates. Loihi 2 [Davies_etal21] which directly processes these events asynchronously in a fast and efficient way. This system effectively preserves privacy as it processes the sparse event stream in real-time which is both important for quick detection but also to discard events as soon as they are processed. In addition, it reduces energy consumption by exploiting the sparsity of input events and neural networks activations which is important for always-on inference, especially when a battery-powered operation is needed. Beyond the fall detection application, this smart security camera is a proof-of-concept of real-time always-on perception systems at the edge where latency, energy consumption, and privacy matter.
@@
In the last five years, event-based vision sensing has been extensively explored to solve many classes of computer vision problems, such as classification, segmentation, detection, and tracking [Gallego_etal22, Cazzato_Bono24, Chen_etal25, Yang_etal25]. However, when targeting a fully embedded system, EVS requires dedicated processing to realize the latency and energy efficiency advantages of the sensor at the system level. In particular, asynchronous near-memory processing using digital neuromorphic chips is a natural fit in terms of processing paradigm and technology maturity to exploit the high temporal resolution and sparsity of event streams [Christensen_etal22, Schuman_etal22, Kudithipudi_et25]. This end-to-end neuromorphic sensing and processing approach was particularly studied using the Intel Loihi research chip [Davies_etal21], which was shown to be 3-30 more energy-efficient than the embedded Nvidia GPUs [Paredes-Valles_etal24, Ceolini_etal20].
@@
The main limitation of these works is the USB-based interface, which allows for high flexibility but imposes bottlenecks on the speed and efficiency of the system. Therefore, there is growing interest for direct interfaces between EVS and embedded processors. For example, dedicated interfaces have been developed to connect EVS with FPGA-based processing [Bonazzi_etal25]. However, these systems remain power-hungry as they target programmability over efficiency. Other systems in research [Rutishauser_etal23] and industry [Yao_etal24, Caccavella_etal24] propose dedicated interfaces on ASICs, but they use low-resolution EVS (typically pixels) and embed very constrained processors, limiting the system’s accuracy and robustness for complex computer vision tasks.
@@ II-A Sony IMX636 sensor
EVS are bio-inspired sensors that asynchronously output per-pixel luminance changes in the form of events. Each pixel independently emits an event when the log-brightness difference since the last event reaches a threshold [Gallego_etal22]. An event is denoted by , where and are the pixel coordinates, is the timestamp with s resolution and is the polarity, i.e., positive or negative for an increase or decrease in light intensity, respectively. This differential sensing paradigm leverages the redundancy in the natural environment, and outputs sparse events with low latency (less than 100 s at 1 KLux), high temporal resolution (s), low power (32 mW at 0.1 M events/s, 73 mW at 300 M events/s) and high dynamic range (beyond 120 dB) [Sony_EVS, Prophesee_IMX636].
The IMX636 sensor [Finateu_etal20] was developed in a collaboration between Sony and Prophesee, combining Sony Semiconductor Solutions’s CMOS image sensor technology with Prophesee’s event-based vision sensing technology. In particular, the light-receiving and luminance detection circuits are incorporated on different layers: a pixel chip with a programmable Region of Interest (ROI), and a logic chip which includes integrated signal processing circuits (e.g., anti-flicker, event filter, and event rate control). These chips are stacked and connected using Cu-Cu technology within each pixel. The IMX636 sensor has an HD resolution of 1280720 pixels, using the industry’s smallest pixel of 4.86 m among stacked EVS, integrated with the logic chip in a 40 nm process [Sony_EVS, Prophesee_IMX636].
@@ II-B Intel Loihi 2 processor
Loihi 2 is the second generation of the Loihi research chip [Davies_etal18, Intel_Loihi2], a fully digital asynchronous chip manufactured using Intel 4 process technology. It uses 128 near-memory asynchronous processing neuro-cores and two networks-on-chip for sparse communication among these cores. It also includes six embedded, synchronous x86 Lakemont cores for management, such as network configuration and data encoding/decoding. Multiple Loihi 2 chips can be combined to form larger systems; Fig. 1 shows a KP system with eight Loihi 2 chips (bottom right).
Loihi 2 uses barrier synchronization between cores to ensure that all neurons compute within the same global algorithmic timestep in a deterministic way. With stored data, the chip operates as fast as possible, whereas with streamed data, the chip can operate based on physical time (using the clock of the x86 cores), event count, or any signal from an external host or device. In addition, algorithmically, Loihi 2 supports sparse (beyond conventional spiking) neural network deployment due to its programmable neuro-core and graded spike message-passing [Shrestha_etal24]. The Loihi 2 inference can therefore be equivalent in accuracy compared to the same quantized model running on an ANN accelerator, and more compute-efficient by exploiting the unstructured sparsity [Zhou_etal25] of synaptic operations (SynOps).
@@ III-B Neural network architectures
MCUNet. MCUNet is a CNN-based model designed to run on tiny hardware resources for always-on inference in IoT devices [Lin_etal20, Lin_etal21]. It has 18 MobileNet inverted residual blocks [Sandler_etal18], characterized by depthwise separable convolution [Brito_etal25], linear bottlenecks, inverted residuals, and non-linear activation. Nevertheless, it is possible to use an arbitrary number of blocks, which is denoted by MCUXB, where X refers to the number of blocks. For instance, we only use 13 blocks to deploy the model on a single Loihi 2 chip. Notably, this is the first time that such a State-Of-The-Art (SOTA) CNN architecture is deployed on neuromorphic hardware.
S4D. On top of the stateless CNN/MCUNet spatial feature extractors and after average pooling and flattening, a State Space Model (SSM) is used for temporal feature extraction. SSMs have recently been positioned as a promising alternative to Transformers for sequence modeling [Patro_Agneeswaran25]. In their convolutional form, unlike recurrent neural networks, they enable highly parallel training on GPUs with stored data. In their recurrent form, they do not suffer from the quadratic scaling of compute cost of the attention mechanism of Transformers, allowing for linear scaling and efficient inference with streaming data. The S4D model [Gu_etal22] is a variant of the Structured State Space for Sequence Modeling (S4) architecture, where a diagonal state space matrix is used to simplify the kernel computation. It was successfully deployed on Loihi 2 [Meyer_etal24].
@@ III-C Neuron models
LIF with binary/graded spikes. LIF neurons are behavioral models of biological neurons, derived from the spike response model [Gerstner_Kistler02]. We use the current-based LIF neuron which is useful for learning patterns with a rich temporal structure [Bouanane_etal23]. The LIF dynamics are expressed as follows:
@@
where is the neuron’s state, , and are the local parameters and is the output. In fact, the recurrent weight matrix , the input projection matrix , and the output projection matrix of the S4D model are diagonalized, so their diagonal values become local dynamical parameters per neuron, analogous to the synaptic and membrane decay factors (or time constants) of LIF neurons [Bal_Sengupta25, Fabre_etal25].
@@ III-D Patched inference
Despite using only 13 blocks of MCUNet, we need around 10 Loihi 2 chips to deploy the entire network due to the large feature map sizes (i.e., number of neurons). However, since the model is stateless, there is no need to keep the neurons states for consecutive time steps. Therefore, we adapt and use the patched inference proposed in [Lin_etal21], where the authors split the input into multiple patches, and process multiple layers per patch to reduce the peak memory requirement of the model. Similarly, we split the 160x160 pixels input into 25 patches of 40x40 pixels as shown in Fig. 3, with a stride of 30 pixels (i.e., overlap of 10 pixels).
The patches are processed sequentially, where each patch goes through the whole 13 blocks of MCU13B (not just a few layers as in [Lin_etal21]) and only the final output feature map is saved. By reusing the Loihi 2 neuro-core memories for each patch, this input-patched inference drastically reduces the memory requirements by an order of magnitude. Once all the patches are processed, the complete output feature map is reconstructed and the S4D inference is triggered. It is to note that the 10 pixels overlap is not enough to produce algorithmically equivalent feature extraction, and this design choice was motivated by the need to fit the model in a single Loihi 2 chip. It is therefore expected to have some loss in feature map quality and overall detection accuracy.
@@ III-E Training
Event representation. The asynchronous events produced by EVS must be accumulated into timesteps for inference on Loihi 2, and for training on GPUs. We accumulate events into constant temporal windows (same as the algorithmic timestep duration) and apply a per-pixel and per-polarity count to get a single graded input spike. The only exception is the binary LIF model where we use binary input spikes (i.e., at least one event was received within the corresponding temporal window) to keep the binary message-passing paradigm.
@@
Focal loss. Our fall detection dataset is highly imbalanced ( of falling actions); therefore, we adopt the Focal Loss (FL) [Lin_etal20_loss] to mitigate class imbalance, defined as follows:
@@
The model’s estimated probability is calculated based on the sum of output spikes over the sample’s duration for the spike-based models, whereas it is calculated based on the maximum logit difference over the sample’s duration for the continuous-output models.
Backpropagation. While ReLU- and S4D-based models are trained using standard Backpropagation (BP), LIF and SigmaDelta spiking models require a custom BP to address the non-differentiability limitation of their activation functions [Neftci_etal19, Eshraghian_etal23]. In this work, we use the SLAYER framework [Shrestha_Orchard18] for direct training of SNNs, as it is widely used and takes into account the hardware constraints (e.g., weight and state quantization) of the Loihi 2 chip.
SLAYER is a form of Backpropagation Through Time (BPTT) which uses surrogate gradients in the backward pass [Neftci_etal19], as shown in Fig. 4. In addition, gradients are computed jointly in time, reducing the complexity of sequential BPTT by parallelized and GPU-accelerated joint gradient computation [Bauer_etal23]. Although it is not mathematically equivalent to BPTT, SLAYER does not differ significantly, in particular when the spiking rate is low which is one of the main goals when using SNNs.
@@ IV Experiments and results
In this section, we present our Loihi 2-aware algorithmic exploration and benchmarking with our newly recorded dataset, to understand the tradeoffs between accuracy and efficiency.
@@ IV-A Fall detection benchmark
Metrics. We follow the benchmarking principles defined in NeuroBench [Yik_etal25], with algorithmic- and system-level evaluations. At the algorithmic level, for a pre-defined throughput of 16 predictions/s, we evaluate in software:
@@ IV-C System-level evaluation
The Loihi 2 deployment and inference was performed on a KP revision C2 board with a N3C1 Loihi2 chip, using the NxKernel 0.2.1 library. We mapped every model trying to minimize the number of Loihi 2 cores, and we verified that the KP-EVS hardware is electrically functional with an event stream passing through the interface into Loihi 2. However, it was not used in the system-level evaluation which is based on stored data, sending recorded events from the supoerhost PC to the KP board using the 10G Ethernet interface.
Latency. First, we need to define the two processing schemes of Loihi 2 [Meyer_etal24]: (1) pipelined processing where each layer is processed once at each hardware timestep, and (2) fall-through processing where the input goes through all the layers first in multiple hardware timesteps before moving to the next input. For the CNN-based models with input-parallel processing, a fall-through processing is used and a maximum throughput of 500 hz can be achieved, with a minimum input-output latency of 2 ms. For the MCUN13B+S4D model with input-patched processing, a pipelined processing of the MCU13B backbone is used with 38 hardware steps (13 blocks + 25 input patches), followed by a fall-through processing of the S4D blocks and output linear layers. A maximum throughput of 25 Hz can be achieved with a minimum input-output latency of 40 ms. Therefore, all models are deployed in a single Loihi 2 chip and satisfy their real-time constraints by effectively reaching (and exceeding) the target 16 predictions/s. In the real-world inference, the Loihi 2 chip is slowed-down to operate in real-time, with an input-output latency of 60 ms which is negligible for the fall detection application that does not require a lower latency response.
@@
| Pareto front model |
Power (mW) | ||||
| Static | Dynamic | Total | ||||
| CNN + MLP [Graded LIF] | 61 | 46.0 | 0.3 | 46.3 | ||
| CNN + S4D | 80 | 76.0 | 1.5 | 77,5 | ||
| MCU13B + S4D | 99 | 80,7 | 8.2 | 88,9 | ||

### v2

@@ I Introduction
The world population is rapidly growing with an increase in life expectancy, particularly in developed countries [5]. Consequently, elderly care is a global concern which is gaining attention. In particular, falls are becoming a major public health problem for the elderly [42]. According to the World Health Organization, falls are the second most common cause of unintentional injury-related deaths worldwide, after road traffic accidents [22]. Therefore, there is a growing need for fall detection systems in the healthcare industry. A fall detection system aims to trigger an alert after a fall has occurred to emergency caregivers [57]. On the one hand, quick response is critical after a fall and medical outcomes of falls as well as the associated healthcare expenses largely depend on rescue time [39, 27]. On the other hand, it has been shown that fall detection systems decrease the fear of falling, which in turn can decrease the likelihood of a fall [24, 39].
@@
Fall detection systems use two main types of sensors [12]: body-worn inertial sensors and vision sensors. Accelerometers and gyroscopes embedded in smartphones and smartwatches measure body motion independently of environmental factors (e.g., objects, lighting). However, continuous wear and regular recharging limit their practicality, particularly for elderly people [2]. In contrast, vision sensors like RGB cameras, infrared cameras, depth cameras, and 3D-based methods using camera arrays are non-intrusive for the user and provide continuous passive monitoring without depending on user compliance. They are therefore useful in settings such as nursing homes, hospitals, or assisted living facilities. Combined with deep learning models, vision-based solutions can be highly accurate and adaptable to different environments [54]. Yet, privacy concerns remain substantial [53, 30, 47] as residents often oppose in-room cameras [58]. Studies on fall detection usually lack strategies to effectively ensure data privacy, which prevents their real-life deployment [30].
@@
In this work, we propose a privacy-preserving vision-based solution that consists of the Sony®111Sony is a registered trademark of Sony Group Corporation or its affiliates. Event-based Vision Sensor (EVS) [23] which only captures per-pixel changes of light intensity in the form of events, combined with the Intel®222Intel is a registered trademark of Intel Corporation or its affiliates. Loihi 2 [19] which directly processes these events asynchronously in a fast and efficient way. This system effectively preserves privacy by processing the sparse event stream in real time, enabling rapid detection while ensuring that events are discarded immediately after processing at the edge. In addition, it reduces energy consumption by exploiting the sparsity of input events and neural networks activations which is important for always-on inference, especially when a battery-powered operation is needed. Beyond the fall detection application, this smart security camera is a proof-of-concept of real-time always-on perception systems at the edge where latency, energy consumption, and privacy matter.
@@
In the last five years, event-based vision sensing has been extensively explored to solve many classes of computer vision problems such as classification, segmentation, detection, and tracking [25, 15, 10, 13, 59]. However, when targeting a fully embedded system, EVS requires dedicated processing to realize the latency and energy efficiency advantages of the sensor at the system level. In particular, asynchronous near-memory processing using digital neuromorphic chips is a natural fit in terms of processing paradigm and technology maturity to exploit the high temporal resolution and sparsity of event streams [14, 52, 32]. This end-to-end neuromorphic sensing and processing approach was particularly studied using the Intel Loihi research chip [19], which was shown to be 3-30 more energy-efficient than the embedded Nvidia GPUs [43, 11].
@@
The main limitation of these works is the USB-based interface, which allows for high flexibility but imposes bottlenecks on the speed and efficiency of the system. Therefore, there is growing interest for direct interfaces between EVS and embedded processors. For example, dedicated interfaces have been developed to connect EVS with FPGA-based processing [6]. However, these systems remain power-hungry as they target programmability over efficiency. Other systems in research [50] and industry [60, 9] propose dedicated interfaces on ASICs, but they use low-resolution EVS (typically pixels) and embed very constrained processors, limiting the system’s accuracy and robustness for complex computer vision tasks.
@@ II-A Sony IMX636 sensor
EVS are bio-inspired sensors that asynchronously output per-pixel luminance changes in the form of events. Each pixel independently emits an event when the log-brightness difference since the last event reaches a threshold [25]. An event is denoted by , where and are the pixel coordinates, is the timestamp with s resolution and is the polarity, i.e., positive or negative for an increase or decrease in light intensity, respectively. This differential sensing paradigm leverages the redundancy in the natural environment, and outputs sparse events with low latency (less than 100 s at 1 KLux), high temporal resolution (s), low power (32 mW at 0.1 M events/s, 73 mW at 300 M events/s) and high dynamic range (beyond 120 dB) [16, 48].
The IMX636 sensor [23] was developed in a collaboration between Sony and Prophesee, combining Sony Semiconductor Solutions’s CMOS image sensor technology with Prophesee’s event-based vision sensing technology. In particular, the light-receiving and luminance detection circuits are incorporated on different layers: a pixel chip with a programmable Region of Interest (ROI), and a logic chip which includes integrated signal processing circuits (e.g., anti-flicker, event filter, and event rate control). These chips are stacked and connected using Cu-Cu technology within each pixel. The IMX636 sensor has an HD resolution of 1280720 pixels, using the industry’s smallest pixel of 4.86 m among stacked EVS, integrated with the logic chip in a 40 nm process [16, 48].
@@ II-B Intel Loihi 2 processor
Loihi 2 is the second generation of the Loihi research chip [18, 33], a fully digital asynchronous chip manufactured using Intel 4 process technology. It uses 128 near-memory asynchronous processing neuro-cores and 2 networks-on-chip for sparse communication among these cores. It also includes 6 embedded, synchronous x86 Lakemont cores for management, such as network configuration and data encoding/decoding. Multiple Loihi 2 chips can be combined to form larger systems; Fig. 1 shows a KP system with 8 Loihi 2 chips.
Loihi 2 uses barrier synchronization between cores to ensure that all neurons compute within the same global algorithmic timestep in a deterministic way [18, 34]. With stored data, the chip operates as fast as possible, whereas with streamed data, the chip can operate based on physical time (using the clock of the x86 cores), event count, or any signal from an external host or device. In addition, algorithmically, Loihi 2 supports sparse (beyond conventional spiking) neural network deployment due to its programmable neuro-core and graded spike message-passing [56]. The Loihi 2 inference can therefore match the accuracy of the same quantized model on an ANN accelerator, while achieving higher computing efficiency by exploiting the unstructured sparsity [63, 62] of synaptic operations (SynOps). Using Loihi 2, we aim not for biological plausibility, but for energy efficiency via biological inspiration.
@@ III-B Neural network architectures
MCUNet. MCUNet is a CNN-based model designed to run on tiny hardware resources for always-on inference in IoT devices [36, 35]. It has 18 MobileNet inverted residual blocks [51], characterized by depthwise separable convolution [8], linear bottlenecks, inverted residuals, and non-linear activation. Nevertheless, it is possible to use an arbitrary number of blocks, which is denoted by MCUXB, where X refers to the number of blocks. For instance, we only use 13 blocks to deploy the model on a single Loihi 2 chip. Notably, this is the first time that such a State-Of-The-Art (SOTA) CNN architecture is deployed on neuromorphic hardware.
S4D. On top of the stateless CNN/MCUNet spatial feature extractors and after average pooling and flattening, a State Space Model (SSM) is used for temporal feature extraction. SSMs have recently been positioned as a promising alternative to Transformers for sequence modeling [46]. In their convolutional form, unlike recurrent neural networks, they enable highly parallel training on GPUs with stored data. In their recurrent form, they do not suffer from the quadratic scaling of compute cost of the attention mechanism of Transformers, allowing for linear scaling and efficient inference with streaming data. The S4D model [29] is a variant of the Structured State Space for Sequence Modeling (S4) architecture, where a diagonal state space matrix is used to simplify the kernel computation. It was successfully deployed on Loihi 2 [38].
@@ III-C Neuron models
LIF with binary/graded spikes. LIF neurons are behavioral models of biological neurons, derived from the spike response model [26]. We use the current-based LIF neuron which is useful for learning patterns with a rich temporal structure [7]. The LIF dynamics are expressed as follows:
@@
where is the neuron’s state, , and are the local parameters and is the output. In fact, the recurrent weight matrix , the input projection matrix , and the output projection matrix of the S4D model are diagonalized, so their diagonal values become local dynamical parameters per neuron, analogous to the synaptic and membrane decay factors (or time constants) of LIF neurons [3, 21].
@@ III-D Patched inference
Despite using only 13 blocks of MCUNet, we need around 10 Loihi 2 chips to deploy the entire network due to the large feature map sizes (i.e., number of neurons). However, since the model is stateless, there is no need to keep the neurons states for consecutive time steps. Therefore, we adapt and use the patched inference proposed in [35], where the authors split the input into multiple patches, and process multiple layers per patch to reduce the peak memory requirement of the model. Similarly, we split the 160x160 pixels input into 25 patches of 40x40 pixels as shown in Fig. 3, with a stride of 30 pixels (i.e., overlap of 10 pixels).
The patches are processed sequentially, where each patch goes through the whole 13 blocks of MCU13B (not just a few layers as in [35]) and only the final output feature map is saved. By reusing the Loihi 2 neuro-core memories for each patch, this input-patched inference drastically reduces the memory requirements by an order of magnitude. Once all the patches are processed, the complete output feature map is reconstructed and the S4D inference is triggered. It is to note that the 10 pixels overlap is not enough to produce algorithmically equivalent feature extraction, and this design choice was motivated by the need to fit the model in a single Loihi 2 chip. It is therefore expected to have some loss in feature map quality and overall detection accuracy.
@@ III-E Training
Events encoding. The asynchronous events produced by EVS must be accumulated into algorithmic timesteps for inference on Loihi 2 and for training on GPUs. We accumulate events into constant temporal windows (i.e., algorithmic timestep duration) and apply a per-pixel and per-polarity count to get a single graded input spike. The only exception is the binary LIF model where we use binary input spikes (i.e., at least one event was received). During Loihi 2 timestep-based inference, event frames are not built, and events are simply encoded into sparse graded spikes.
@@
Focal loss. Our fall detection dataset is highly imbalanced ( of falling actions); therefore, we adopt the Focal Loss (FL) [37] to mitigate class imbalance, defined as follows:
@@
The model’s estimated probability is calculated over the sample’s duration based on the sum of output spikes for the spike-based models, whereas it is based on the maximum logit difference the continuous-output models.
Backpropagation. While ReLU- and S4D-based models are trained using standard Backpropagation (BP), LIF and SigmaDelta spiking models require a custom BP to address the non-differentiability limitation of their activation functions [41, 20]. In this work, we use the SLAYER framework [55] for direct training of SNNs, as it is widely used and takes into account the hardware constraints (e.g., weight and state quantization) of the Loihi 2 chip.
SLAYER is a form of Backpropagation Through Time (BPTT) which uses surrogate gradients in the backward pass [41], as shown in Fig. 4. In addition, gradients are computed jointly in time, reducing the complexity of sequential BPTT by parallelized and GPU-accelerated joint gradient computation [4]. Although it is not mathematically equivalent to BPTT, SLAYER does not differ significantly, in particular when the spiking rate is low.
@@ IV Experiments and results
In this section, we present our Loihi 2-aware algorithmic exploration and benchmarking with our newly recorded dataset, highlighting the tradeoffs between accuracy and efficiency.
@@ IV-A Fall detection benchmark
Metrics. We follow the benchmarking principles defined in NeuroBench [61], with algorithmic- and system-level evaluations. At the algorithmic level, for a pre-defined throughput of 16 predictions/s, we evaluate in software:
@@ IV-C System-level evaluation
The Loihi 2 deployment and inference was performed on a KP revision C2 board with a N3C1 Loihi2 chip, using the NxKernel 0.2.1 library. We mapped every model trying to minimize the number of Loihi 2 cores, and we verified that the KP-EVS hardware is electrically functional with an event stream passing through the interface into Loihi 2. However, it was not used in the system-level evaluation which is based on stored data, sending recorded events from the superhost PC to the KP board using the 10G Ethernet interface.
Latency. First, we need to define the two processing schemes of Loihi 2 [38]: (1) pipelined processing where each layer is processed once at each hardware timestep, and (2) fall-through processing where the input goes through all the layers first in multiple hardware timesteps before moving to the next input. For the CNN-based models with input-parallel processing, a fall-through processing is used and a maximum throughput of 500 hz can be achieved, with a minimum input-output latency of 2 ms. For the MCUN13B+S4D model with input-patched processing, a pipelined processing of the MCU13B backbone is used with 38 hardware steps (13 blocks + 25 input patches), followed by a fall-through processing of the S4D blocks and output linear layers. A maximum throughput of 25 Hz can be achieved with a minimum input-output latency of 40 ms. Therefore, all models are deployed in a single Loihi 2 chip and satisfy their real-time constraints by effectively reaching (and exceeding) the target 16 predictions/s. In the real-world inference, the Loihi 2 chip is slowed-down to operate in real-time, with an input-output latency of 60 ms which is negligible for the fall detection application that does not require a lower latency response.
@@
| Pareto front model | Loihi 2 cores | Power (mW) | ||
| Static | Dynamic | Total | ||
| CNN + MLP [Graded LIF] | 61 | 46.0 | 0.3 | 46.3 |
| CNN + S4D | 80 | 76.0 | 1.5 | 77,5 |
| MCU13B + S4D | 99 | 80,7 | 8.2 | 88,9 |